the sum of two numbers is 30 and the difference is 6

Answer by ankor@dixie-net.com(22740)

the sum of two numbers is 30
a + b = 30
and the difference is 6
a - b = 6
:
Find a using elimination
a + b = 30
a - b = 6
-------------adding eliminates b, find a
2a + 0 = 36
a = 36/2
a = 18
Use the 2nd equation to find b
18 - b = 6
b = 12 obviously
:
:
Check solutions in the 1st equation
18 + 12 = 30